You are considering the purchase of 1000 ordinary shares in CBA Bank Ltd. CBA is a commercial bank predominantly operating in Australia and New Zealand. CBA only pays a dividend once a year to shareholders but has a very generous dividend payout ratio of 70%. Annual Earnings per Share are currently A$7.50. Security analysts have a consensus view that CBA will succeed in growing these earnings at a rate of 3% per year


The current market risk premium is 6% from investing in major listed shares in Australia. CBA currently has a beta of 1.2. This is worrying you as you have done some research and find that the average historical beta for CBA over the last 20 years has been 1.0. A recent graduate of MIT who studied this unit has told you that the high beta is because of enhanced volatility around CBA returns stemming from the fact that it has been involved in several recent scandals. You find the following article on one of those incidents

“The CBA share price has been slammed after the government regulator AUSTRAC (Australian Transactions Reports & Analysis Centre) accused it of failing to enforce checks designed to stop criminals or tax dodgers depositing cash in the bank (commonly via anonymous “Intelligent Deposit Machines”) in order to hide their ill-gotten gains. AUSTRAC is seeking “orders for civil pecuniary penalties” against the CBA and, given the serious allegations, CBA could face an enormous fine”

You look at the CBA share price on the internet and see it is trading currently at $75. It was trading at $81 before the AUSTRAC allegations were made public.

You are of the opinion that CBA will be able to maintain a 70% dividend pay-out ratio even in the event of having to pay a high level of penalties in relation to the allegations.

(a) If the yield on 10 year Australian government bonds is currently 2.5%, what will your required rate of return be as a CBA ordinary shareholder? Show all workings

(b) Should you buy the shares at their current trading price? You need to run a calculation as well as provide reasoning.
Answer #1

Solution :

Risk-free rate = 2.5 %, Market risk premium = 6 %, Beta = 1.2 ( We will consider this beta as opposed to 1.00 because due to certain issues systematic risk has increased )

Using CAPM we can calculate the cost of equity

Cost of equity = Risk free rate + Beta * risk premium = 2.5% + 1.2 * 6% = 9.7%

Part B ) We have EPS of 7.5 and dividend payout is 70% . Dividend (0)= 70% * 7.5 = 5.25

growth rate = 3% forever , Dividend (1) = 5.25 * ( 1 +growth ) = 5.25 * 1.03 = 5.4075

Cost of equity = 9.7%

We would used Dividend discount model to find the current price

Price = D 1 / ( cost of equity - growth ) = 5.4075 / (0.097-0.03) = 5.4075 / 0.67 = 80.71

So justifiable value according to DDM model is 80.71 and current price is 75. So, You can consider buying it

We have taken beta as 1.2 and earlier beta was 1.00 so if we reduce the value of beta then cost of equity will go down. And theoretical share price will be higher than 80.71.

This 80.71 is conservative price and still more than current trading price of 75.